OpenStudy (anonymous):

Calculate the volume of a cylinder with radius of 2 ft. and height of 3 ft

OpenStudy (anonymous):

Volume of a cylinder = π(r^2)(h). You've got r=2 & h=3.

OpenStudy (anonymous):

Um what does this thing mean? -> ^

OpenStudy (anonymous):

Ah sorry. I'll rewrite it so it's clearer:\[πr^2h\] That means you square the number, i.e. multiply it by itself. So r^2 would be 2^2 = 2 * 2 = 4. And r^3 would be 2 * 2 * 2 = 8.

OpenStudy (anonymous):

I got 37.68?

OpenStudy (anonymous):

Yep, that's right!